A cup of coffee is poured at 190˚F and is allowed to cool in a 70˚F room. If the temperature of the coffee is 170˚F after half an hour,
(a) What will the temperature be after 70 minutes?
(b)How long it will take the coffee to cool to 120 ˚F?

Found 2 solutions by CubeyThePenguin, MathTherapy:
Answer by CubeyThePenguin(3113)

190 F after 0 minutes, 170 F after 30 minutes --> decrease of 2/3 degrees every minute

a) 190 - 70(2/3) = 430/3 = 143 1/3 degrees

b) 190 - (2/3)x = 120
70 = 2x/3
210 = 2x

x = 105 minutes

Answer by MathTherapy(10556)

A cup of coffee is poured at 190˚F and is allowed to cool in a 70˚F room. If the temperature of the coffee is 170˚F after half an hour,
(a) What will the temperature be after 70 minutes?
(b)How long it will take the coffee to cool to 120 ˚F?
The other person, CUBEYthePENGUIN is WRONG!
You can use Newton's Law of Cooling to get a temperature of:
In addition, to reach a temperature of 120oF will take